OpenStudy (anonymous):

what is sec^2 theta-tan^2theta=1 prove it

OpenStudy (anonymous):

well lets switch this into sin and cos \[\frac{ 1 }{ \cos^2\theta } - \frac{ \sin^2\theta }{\cos^2\theta } = 1\] which becomes \[\frac{ 1 - \sin^2\theta }{ \cos^2\theta } = 1\] Now knowing that there is a trig property where \[\sin^2\theta = \cos^2\theta = 1\] Solve that for cos^2theta \[1-\sin^2\theta = \cos^2\theta\] So this means in the original that the 1-sin^2theta = cos^2theta so you can replace it with that \[\frac{ \cos^2\theta }{ \cos^2\theta } = 1\] \[1=1\] This is your proof

OpenStudy (anonymous):

should be sin^2theta + cos^2theta = 1 in middle there